OpenStudy (anonymous):

A little help please about integral Hey guys , today we took some examples on integrals , but there's something wrong with the result of this one, either I didn't focus when I was writing or when I tried to solve it myself my result was wrong :D I'll post it + my solution in a sec

OpenStudy (anonymous):

the integral: \[\sin^2x \cos^2x dx\] my final solution: \[\frac{ 1 }{ 8} x -\frac{ 1 }{ 32 }*\sin^4x +c\]

OpenStudy (anonymous):